OpenStudy (anonymous):

Which one of the following is NOT an example of work being done? The moon orbiting Earth Pushing a box from the bottom of a hill to the top of the hill Pulling a sled across a field covered with snow Lifting a bookbag off the floor

OpenStudy (anonymous):

If the path of moon around the earth is assumed to be circular, the direction of the centripetal force of the earth and the direction of displacement of moon at any point are perpendicular to each other W=F.s W=Fscos@ @=90 Therefore Work done is 0

OpenStudy (anonymous):

so it was the moon orbiting earth?

OpenStudy (anonymous):

help anyone? is the moon orbiting earth is correct?

OpenStudy (mos1635):

yeap!!

OpenStudy (anonymous):

help please

OpenStudy (experimentx):

according to definition of work done. W = F (dot) R (where R is radial vector) = FRcos\(\theta\) which says that for work to be done, Force and Distance traveled have to be in same direction. But in this case, \(\theta\) is 90 so no work is done.

OpenStudy (anonymous):

so it was the moon orbit the earth right

OpenStudy (experimentx):

yes 99.99% right.

OpenStudy (mos1635):

100.01%

OpenStudy (anonymous):

thanks everyone

OpenStudy (mos1635):

and then some :)

OpenStudy (experimentx):

on average 100% sure.

OpenStudy (mos1635):

thumbs up